a small company invested $20,500 such that a part of the money is in an account with a 4% invest rate and the rest at a 5% rate. the annual interest from the 5% account is $35 more than the interest earned from the 4% account. find the account of money the company invested at each rate.
Answer by lwsshak3(11628)

a small company invested $20,500 such that a part of the money is in an account with a 4% invest rate and the rest at a 5% rate. the annual interest from the 5% account is $35 more than the interest earned from the 4% account. find the account of money the company invested at each rate.
***
let x=amt invested in the 4% account
20500-x=amt invested in the 5% account
..
5%(20500-x)-4%x=35
1025-.05x-.04x=35
.09x=990
x=11000
20500-x=9500
amt invested in the 4% account=$11,000
amt invested in the 5% account=$9,500